ADNAN MUHAMMED

Grade 12,

A man is weighing 160 pounds stans on the floor of elevator. What force will be exerted on the floor what if it 1- decends With a uniform acceleration of 1 ft/sec2 2- Ascending with a uniform acceleration of 1ft/sec2.

A man is weighing 160 pounds stans on the floor of elevator. What force will be exerted on the floor what if it 1- decends With a uniform acceleration of 1 ft/sec2
2-  Ascending with a uniform acceleration of 1ft/sec2.

Grade:12th pass

4 Answers

Pooja
4 years ago
g = 32.2ft /s2
force exerted on the floor =N
➡️when lift decends
N = m(g-a)
   =160(32.2-1)
   = 4976 pound.ft/sec2
➡️when lift ascends
N = m(g+a)
     =160(32.2+1)
     = 5132 pound.ft/sec2
yohib ahmed
4 years ago
g = 32.2ft /s2??
 
please ask me why g = 32.2ft /s2 you put here?
 
please ask me why g = 32.2ft /s2 you put here?
yohib ahmed
4 years ago
 
g = 32.2ft /s2??
 
please ask me why g = 32.2ft /s2 you put here?
 
please ask me why g = 32.2ft /s2 you put here?
 
 
I got it. You convert m into ft 9.8x3.28 
Khimraj
4 years ago
 
when lift decends
N = m(g-a)
   =160(32.2-1)
   = 4976 pound.ft/sec2
when lift ascends
N = m(g+a)
     =160(32.2+1)
     = 5132 pound.ft/sec2